the equation x^2+y^2=25 represents a______.
Found 2 solutions by Alan3354, DrBeeee:
Answer by Alan3354(69443)

the equation x^2+y^2=25 represents
------------
Hint:
The x^2 and y^2 terms have the same sign and the same coefficient.

Answer by DrBeeee(684)

The general equation of a circle is
(x-x0)^2+(y-y0)^2 = r^2 , where (x0,y0) are the coordinates of the circle's origin and r is the radius of the circle.
Comparing your equation to the general equation of a circle leads to the conclusion that your equation is that of:
A circle with center at the origin and radius of 5 units.